cooliopranks802:

Simplify 27 to the negative 2 thirds power.

Eiwoh2:

You could probably do this on an algebra calculator. If not, do you know the first step in simplyfying equations like these?

dude:

Since you're gone, \(27^{-\frac{2}{3}}\) Negative powers will result in a reciprocal \(\frac{1}{27^{\frac{2}{3}}}\) The numerator is the power of the number and the denominator denotes the root ============================ \[\frac{1}{\sqrt[3]{27^2}}\]

Eiwoh2:

2 smart for mi. Also, you're almost 9 9. Congrats, m8. xD